OpenStudy (anonymous):

How do I know the domain of 50 tan theta - 50 theta?

OpenStudy (anonymous):

The 50's don't matter. tan = sin/cos, which only has existence problems when cos = 0, so the domain is all real theta, except for when cos(theta) evaluates to zero, which is when \[\theta \neq \pm n \frac{\pi}{2}\] where n is an odd integer

OpenStudy (anonymous):

typo: the not equal should be equal

OpenStudy (anonymous):

Why is it only values of the first quadrant? I inferred that since you only mentioned pi/2 and not 3pi/2 as well

OpenStudy (anonymous):

Also, tan of pi/2 is undefined because sin is 1 and cos is 0

OpenStudy (cwrw238):

its not only values in first quadrant theta = +/- n pi/2 as uhhh as said